OpenStudy (firejay5):

Simplify. 28. (a^3b^3) x (ab)^-2

OpenStudy (anonymous):

a^3b^3 x a^-2b^-2 = ab

OpenStudy (firejay5):

\[\frac{ (a^3b^3) }{ ab^-2 }\]

OpenStudy (firejay5):

It's multiplying

OpenStudy (anonymous):

(a^b^3)/(ab)^2 = still ab